You are trying to dissolve 11g of glucose in 100 mL water. You then proceeded to dilute a 20 mL sample of that solution to 0.500L water. Solve for the grams of glucose that are in a 100mL of the final solution.



1
Expert's answer
2021-02-03T07:24:33-0500

Solution:

"density = \dfrac{11g}{100ml}=0.11(\frac{g}{ml})"


"m=V*d=20ml*0.11(\frac{g}{ml})=2.2g"

 

"0.500L=500ml"


"m=\dfrac{2.2g}{500ml}*100ml=0.44g"


Answer:

We need to add 0.44 g of glucose to 100 ml of the final solution.